OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago. A. 3(t + 7) = t – 19 B. t + 19 = 3(t – 7) C. t + 7 = 3(t – 19) D. 3(t + 19) = t – 7

Parth (parthkohli):

What was Tim's age 19 years back, if it is \(t\) at the moment?

OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago. A. 3(t + 7) = t – 19 B. t + 19 = 3(t – 7) C. t + 7 = 3(t – 19) D. 3(t + 19) = t – 7

Parth (parthkohli):

Do you know what Tim's age was \(19\) years back if he is \(t\) years old at the moment?

OpenStudy (anonymous):

no sorry

Parth (parthkohli):

What is your age right now?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

can you helpme with the question?

Parth (parthkohli):

What was your age 5 years back?